OpenStudy (anonymous):

limit of an = (2n^2)/(n^2+6) ?

OpenStudy (anonymous):

\[ \n^2+6\approx n^2 \] near infinity So how much is the limit?

OpenStudy (anonymous):

\[ n^2+6 \approx n^2 \]

OpenStudy (anonymous):

So \[ \frac {2 n^2}{n^2+6}\approx \frac {2 n^2}{n^2}=2 \] the limit is 2

OpenStudy (anonymous):

Another way is to do the following \[ \frac{2 n^2}{n^2\left( 1 +\frac 6 {n^2} \right ) }=\frac{2 }{\left( 1 +\frac 6 {n^2} \right ) } \] It is easy now to see that the limit is 2